OpenStudy (anonymous):

Find the first three iterates of the given function for the given initial value. (value in comments) Please Help

OpenStudy (anonymous):

\[f(x)=\frac{ 2 }{x ^{2} }, x _{0}=-\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

¯\(°_o)/¯

OpenStudy (anonymous):

yes I know

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

When they say "Find the first three iterates" they mean "find the first three terms of the recursive sequence"

jimthompson5910 (jim_thompson5910):

Actually, according to this page http://www.glencoe.com/sites/common_assets/support_pages/sc_technologies_2g.pdf it's the first three after the given starting term

jimthompson5910 (jim_thompson5910):

so we'll go with that

jimthompson5910 (jim_thompson5910):

To find x1 (the first iterate), you plug in x = x0 into the given function and simplify

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so what do you get when you plug in x = -1/2

OpenStudy (anonymous):

 If that makes any sense lol.